# Curriculum effects and compositionality emerge with in-context learning in neural networks

### v3

@@ B.3 Concurrent ICL and IWL reproduce both curriculum effects on the category task
Our main hypothesis was that a single network capable of both ICL and IWL would reproduce the interaction between Rotation and Curriculum observed in humans in the category-learning setting [5], with ICL producing a blocking advantage in the Rule-like task and IWL producing an interleaving advantage in the Rotated task. We evaluated metalearned networks’ few-shot test accuracy on the Rule-like task (see Figure 2g in the main text) and their finetuned train accuracy on the Rotated task (see Figure 2h in the main text), and performed the same GLM analysis (see Table 3). We observed significant main effects of both Rotation ($\chi^{2}=52.23$, $p=2.96e-13$) and Curriculum ($\chi^{2}=60.58$, $p=7.07e-15$), as well as a significant interaction ($\chi^{2}=52.19$, $p=5.03e-13$). Analyses of the simple main effects revealed a significant blocking advantage in the Rule-like task (same as ICL result above; $\chi^{2}=5.51$, $p=0.019$), and a significant interleaving advantage in the Rotated task ($\chi^{2}=911$, $p<2.2e-16$). These analyses confirmed our hypothesis that it is possible for a single neural network model to reproduce the blocking advantage observed in humans on the Rule-like task (due to ICL) and the interleaving advantage observed in humans on the Rotated task (due to IWL).
@@ C.2 Metalearning on rotated task
What we have called the Rule-like versions of our tasks are only “rule-like” in the sense that the rules governing the task are simpler to discover because the features determining the correct answers (e.g., line orientation and length, or color and animal identities) are intuitive to humans. The rotated versions of the two tasks are more difficult because the rules of the task are only succinctly describable in rotated feature spaces that are unintuitive. Likewise, the rule-like versions of the tasks are easier for the metalearning networks to solve in-context because the appropriate features are familiar from the task distributions on which they metalearned. If instead the networks were more familiar with these rotated feature spaces, we would expect in-context learning to be easier on tasks whose rules utilized similar feature spaces. To test this hypothesis, we trained metalearning networks on a distribution of rotated tasks in the compositional setting (see Figure 9). As expected, the network showed the opposite pattern of curriculum effects, exhibiting a blocking advantage in the rotated task rather than the rule-like task. The network also suffered from catastrophic forgetting in the Rule-like task rather than the Rotated task. Interestingly, catastrophic forgetting was not as severe compared to the network trained on a distribution of rule-like tasks and finetuned on a rotated task. This meant that by the fourth block, the interleaving advantage observed in through earlier blocks had disappeared. This may suggest that metalearning on the rotated version of the task equipped the network with an inductive bias to learn strategies that suffer from less interference when related trials are blocked over time.
@@
In general, this experiment is consistent with our hypothesis that the blocking advantage should emerge on tasks where the properties of ICL dominate, and the interleaving advantage should emerge on tasks where properties of IWL dominate due to catastrophic forgetting. However, the finetuning results suggest that catastrophic forgetting in IWL can be more or less severe depending on the inductive biases imparted by metalearning.
@@
Figure 9: Results for a metalearning network trained on a distribution of rotated tasks. This network acquires the ability to solve new tasks in-context, but shows the opposite curriculum effects as the networks trained on a distribution of rule-like tasks (compare to Figure 3g-i in the main text). In the few-shot evaluation, the network shows good performance on the rotated version of the task (red) and poor performance on the rule-like task (blue). Importantly, a blocking advantage was observed in the rotated, rather than the rule-like setting, consistent with our hypothesis that the blocking advantage emerges on tasks where ICL is possible. Finetuning results show that IWL promotes catastrophic forgetting on the rule-like version of the task, rather than the rotated task. Interestingly, this catastrophic forgetting was not as severe compared to the network trained on the rule-like task and finetuned on the rotated task (see Figure 3i in the main text), suggesting that the network was able to learn a strategy in the first block that did not interfere as much with the learning in the second block.
C.3 Blocking over context vs. gradient steps
@@
For completeness and to facilitate understanding of our main results, we also performed experiments where these two notions of time are incongruent (i.e., blocked over the context and interleaved over the gradient steps, or interleaved over the context and blocked over the gradient steps). Here, we report such results from the same metalearning model trained on the compositional task (see Figure 3g-i in the main text). After metalearning, the model was finetuned in the usual way but with these extra two incongruent curriculum conditions.
Figure 10: Results of experiments investigating effects of blocking over the study examples given in the context vs. blocking over the gradient steps during finetuning. The finetuning results from Figure 3h-i in the main text report the main cases of interest where trials were either blocked over both context and gradient steps or interleaved over both context and gradient steps. Here, we show these same results alongside the cases where the two are mismatched. When trials are blocked over the context but interleaved over the steps (second row), the model shows good compositional generalization in the rule-like task and does not suffer from catastrophic forgetting. When trials are interleaved over the context but blocked over the steps (third row), the model shows poor compositional generalization and suffers from catastrophic forgetting. These results demonstrate how this model generalizes compositionally only when trials are blocked over the context in the rule-like task, and how catastrophic forgetting occurs whenever large losses are incurred (inducing IWL) and trials are blocked over the gradient steps.
@@
The results are presented alongside the original (congruent) results in Figure 10. In general, the results show that the key dynamics motivating our original theoretical framework hold across these incongruent conditions. In the metalearning model, ICL is successful when encountering new tasks that have familiar structure (in this case, rule-like tasks where related items are blocked over the context). This is shown by the high train (purple and orange) and test (green) accuracy, and the low loss (black) from the beginning of finetuning in the rule-like task when items are blocked over the context (middle column, top two plots). When ICL is successful, compositional generalization performance is good, and little loss is incurred and backpropagated, resulting in less IWL. This can be seen by the lack of catastrophic forgetting in these same two plots. On the other hand, when ICL is unsuccessful (in this case, when the task is rotated or when items are interleaved over the context), compositional generalization is poor, and large losses are backpropagated, resulting in increased IWL. This can be seen in the loss curves (black) in the rotated task (right column), and in the rule-like task when items are interleaved over the context (middle column, bottom two plots). This increased IWL results in catastrophic forgetting when related trials are blocked over the gradient steps. This can be seen in the sharp drops in accuracy on the items trained in the previous block (e.g., TrainA accuracy, shown in purple, drops during the second block as TrainB items are learned) in the rule-like task when items are interleaved over the context but blocked over the steps (middle column, third plot) or in the rotated task when items are blocked over the gradient steps (right column, first and third plots). When trials are randomly interleaved over the gradient steps, no catastrophic forgetting occurs, even in cases where ICL is unsuccessful. This can be seen in the rule-like task when trials are interleaved over both the context and the gradient steps (middle column, fourth plot), and in the rotated task when items are interleaved over the gradient steps (right column, second and fourth plots).

### v4

@@ B.3 Concurrent ICL and IWL reproduce both curriculum effects on the category task
Our main hypothesis was that a single network capable of both ICL and IWL would reproduce the interaction between Rotation and Curriculum observed in humans in the category-learning setting [5], with ICL producing a blocking advantage in the Rule-like task and IWL producing an interleaving advantage in the Rotated task. We evaluated metalearned networks’ few-shot test accuracy on the Rule-like task (see Figure 2g in the main text) and their train accuracy after task-specific training on the Rotated task (see Figure 2h in the main text), and performed the same GLM analysis (see Table 3). We observed significant main effects of both Rotation ($\chi^{2}=52.23$, $p=2.96e-13$) and Curriculum ($\chi^{2}=60.58$, $p=7.07e-15$), as well as a significant interaction ($\chi^{2}=52.19$, $p=5.03e-13$). Analyses of the simple main effects revealed a significant blocking advantage in the Rule-like task (same as ICL result above; $\chi^{2}=5.51$, $p=0.019$), and a significant interleaving advantage in the Rotated task ($\chi^{2}=911$, $p<2.2e-16$). These analyses confirmed our hypothesis that it is possible for a single neural network model to reproduce the blocking advantage observed in humans on the Rule-like task (due to ICL) and the interleaving advantage observed in humans on the Rotated task (due to IWL).
@@ C.2 Metalearning on rotated task
What we have called the Rule-like versions of our tasks are only “rule-like” in the sense that the rules governing the task are simpler to discover because the features determining the correct answers (e.g., line orientation and length, or color and animal identities) are intuitive to humans. The rotated versions of the two tasks are more difficult because the rules of the task are only succinctly describable in rotated feature spaces that are unintuitive. Likewise, the rule-like versions of the tasks are easier for the metalearning networks to solve in-context because the appropriate features are familiar from the task distributions on which they metalearned. If instead the networks were more familiar with these rotated feature spaces, we would expect in-context learning to be easier on tasks whose rules utilized similar feature spaces. To test this hypothesis, we trained metalearning networks on a distribution of rotated tasks in the compositional setting (see Figure 10). As expected, the network showed the opposite pattern of curriculum effects, exhibiting a blocking advantage in the rotated task rather than the rule-like task. The network also suffered from catastrophic forgetting in the Rule-like task rather than the Rotated task. Interestingly, catastrophic forgetting was not as severe compared to the network trained on a distribution of rule-like tasks before undergoing task-specific training on a rotated task. This meant that by the fourth block, the interleaving advantage observed through earlier blocks had disappeared. This may suggest that metalearning on the rotated version of the task equipped the network with an inductive bias to learn strategies that suffer from less interference when related trials are blocked over time.
@@
In general, this experiment is consistent with our hypothesis that the blocking advantage should emerge on tasks where the properties of ICL dominate, and the interleaving advantage should emerge on tasks where properties of IWL dominate due to catastrophic forgetting. However, the task-specific training results suggest that catastrophic forgetting in IWL can be more or less severe depending on the inductive biases imparted by metalearning.
@@
Figure 10: Results for a metalearning network trained on a distribution of rotated tasks. This network acquires the ability to solve new tasks in-context, but shows the opposite curriculum effects as the networks trained on a distribution of rule-like tasks (compare to Figure 3g-i in the main text). In the few-shot evaluation, the network shows good performance on the rotated version of the task (red) and poor performance on the rule-like task (blue). Importantly, a blocking advantage was observed in the rotated, rather than the rule-like setting, consistent with our hypothesis that the blocking advantage emerges on tasks where ICL is possible. Task-specific training results show that IWL promotes catastrophic forgetting on the rule-like version of the task, rather than the rotated task. Interestingly, this catastrophic forgetting was not as severe compared to the network that saw rule-like tasks during metalearning and a rotated task during subsequent task-specific training (see Figure 3i in the main text), suggesting that the network was able to learn a strategy in the first block that did not interfere as much with the learning in the second block.
C.3 Metalearning on interleaved curricula
Our main results show that for both the category-learning task and the compositional task, curriculum effects can depend on whether ICL or IWL is dominant during learning: an interleaving advantage emerged whenever IWL was dominant (e.g., in the rotated versions of the tasks), and a blocking advantage emerged whenever ICL was dominant (e.g., in the rule-like versions of the tasks). While the interleaving advantage is an inherent property of IWL, the blocking advantage we observed in ICL emerged due to the use of blocked curricula during metalearning. For the sake of completeness, here we report results for the compositional task in the case where interleaved curricula were used during metalearning. One might predict that metalearning from interleaved curricula would impart an interleaving advantage, rather than a blocking advantage, to the resulting ICL algorithm. However, it is important to note that if interleaving is defined as random shuffling, a blocked curriculum is a special case of an interleaved curriculum, where the individual examples happen to be organized into blocks. This means that if a model has metalearned to perform well on all possible interleaved curricula, it will have learned an ICL algorithm that is robust to any changes in trial order, and should perform well even when trials are blocked. Figure 11 shows the results from a model after metalearning on interleaved curricula. As expected, the model is capable of learning new tasks via ICL in both the blocked and interleaved curriculum conditions. These results suggest that in our setup, a blocking advantage is not an inherent property of ICL per se, but emerges due to the data distribution used during metalearning.
Figure 11: Results for a metalearning network trained on a distribution of rule-like tasks with interleaved curricula. Due to the asymmetric nature of the blocking vs. interleaving distinction (a blocked curriculum can be seen as a special case of a randomly shuffled, or interleaved, curriculum) this network is capable of learning new tasks via ICL in either curriculum condition.
C.4 Blocking over context vs. gradient steps
@@ C.3 Metalearning on interleaved curricula
For completeness and to facilitate understanding of our main results, we also performed experiments where these two notions of time are incongruent (i.e., blocked over the context and interleaved over the gradient steps, or interleaved over the context and blocked over the gradient steps). Here, we report such results from the same metalearning model trained on the compositional task (see Figure 3g-i in the main text). After metalearning, the model underwent task-specific training in the usual way but with these extra two incongruent curriculum conditions.
@@
The results are presented alongside the original (congruent) results in Figure 12. In general, the results show that the key dynamics motivating our original theoretical framework hold across these incongruent conditions. In the metalearning model, ICL is successful when encountering new tasks that have familiar structure (in this case, rule-like tasks where related items are blocked over the context). This is shown by the high train (purple and orange) and test (green) accuracy, and the low loss (black) from the beginning of task-specific training in the rule-like task when items are blocked over the context (middle column, top two plots). When ICL is successful, compositional generalization performance is good, and little loss is incurred and backpropagated, resulting in less IWL. This can be seen by the lack of catastrophic forgetting in these same two plots. On the other hand, when ICL is unsuccessful (in this case, when the task is rotated or when items are interleaved over the context), compositional generalization is poor, and large losses are backpropagated, resulting in increased IWL. This can be seen in the loss curves (black) in the rotated task (right column), and in the rule-like task when items are interleaved over the context (middle column, bottom two plots). This increased IWL results in catastrophic forgetting when related trials are blocked over the gradient steps. This can be seen in the sharp drops in accuracy on the items trained in the previous block (e.g., TrainA accuracy, shown in purple, drops during the second block as TrainB items are learned) in the rule-like task when items are interleaved over the context but blocked over the steps (middle column, third plot) or in the rotated task when items are blocked over the gradient steps (right column, first and third plots). When trials are randomly interleaved over the gradient steps, no catastrophic forgetting occurs, even in cases where ICL is unsuccessful. This can be seen in the rule-like task when trials are interleaved over both the context and the gradient steps (middle column, fourth plot), and in the rotated task when items are interleaved over the gradient steps (right column, second and fourth plots).
@@
Figure 12: Results of experiments investigating effects of blocking over the study examples given in the context vs. blocking over the gradient steps during task-specific training. The task-specific training results from Figure 3h-i in the main text report the main cases of interest where trials were either blocked over both context and gradient steps or interleaved over both context and gradient steps. Here, we show these same results alongside the cases where the two are mismatched. When trials are blocked over the context but interleaved over the steps (second row), the model shows good compositional generalization in the rule-like task and does not suffer from catastrophic forgetting. When trials are interleaved over the context but blocked over the steps (third row), the model shows poor compositional generalization and suffers from catastrophic forgetting. These results demonstrate how this model generalizes compositionally only when trials are blocked over the context in the rule-like task, and how catastrophic forgetting occurs whenever large losses are incurred (inducing IWL) and trials are blocked over the gradient steps.
C.5 Tradeoff between flexibility and retention in category-learning task
In the main simulations investigating the tradeoff between flexibility and retention (see Figure 5 of the main text), we used the compositional task. Here, we report the results of performing the same manipulation in the category-learning task (see Figure 13). Ablating attention during task-specific training had similar effects in the category-learning setting: greater ablations (larger $p_{a}$) resulted in worse few-shot generalization performance via ICL (left), and required more incremental learning via IWL for models to reach optimal performance (middle left). This meant that more errors were accumulated throughout task-specific training (middle right), resulting in more IWL and better subsequent retention (right).
These results qualitatively reproduce the effects we observed on the compositional task (see Figure 5 of the main text), and are consistent with studies showing a similar tradeoff in human reinforcement learning [9, 17].
Figure 13: Results of ablating attention in category-learning task. Greater ablation ($p_{a}$) resulted in worse few-shot generalization via ICL (left). When the model was not capable of solving the task in context, incremental learning via IWL took longer to reach optimal performance (middle left), and more errors were accumulated over the course of learning (middle right). These errors resulted in greater IWL and better retention when no contextual information was available (right). $p_{a}$ is the probability of inhibiting attention to each example in the context.
C.6 Ablating attention by applying noise to value vectors
The ablation simulations reported above and in the main text (see Figure 5) masked attention to individual examples in the context with probability $p_{a}$. We also experimented with a different kind of ablation, adding Gaussian noise to the attention mechanism itself. This allowed us to replicate our previous ablation results while manipulating the model’s ability to access to contextual information in a different way.
Figure 14: Results of injecting noise into value vectors. Gaussian noise with standard deviation $\sigma_{v}$ was added to the value vectors across all tokens in the context. In both tasks, injecting noise in this way had effects similar to those observed when attention was masked, revealing the same tradeoff between flexibility and retention. More noise (high $\sigma_{v}$) resulted in worse flexibility (few-shot generalization via ICL; left), slower incremental learning via IWL (middle left), more errors (middle right), and greater retention (right).
We added Gaussian noise with standard deviation $\sigma_{v}$ to the value vectors corresponding to all tokens in the context. This was accomplished by running a forward pass on the model, collecting the past keys and values, and applying the noise to the value vectors before running the forward pass again to produce predictions. We used the same experimental design as was used in the simulations that ablated attention using a mask (see above). Again, models that metalearned on unrotated tasks with interleaved curricula were finetuned under different amounts of noise. Few-shot generalization performance was measured with different levels of noise applied at inference time. Different levels of noise were applied throughout finetuning, while recording training accuracy (incremental learning via IWL) and errors (i.e., loss). Retention was tested by applying the maximum amount of noise ($\sigma_{v}=8.0$) at inference time after finetuning.
The results are shown in Figure 14. Again, we observed the same tradeoff between flexibility and retention exhibited by humans [9, 17]. In both tasks, noise degraded ICL few-shot generalization performance, but improved later retention. These results replicate our original findings using a different manipulation, and highlight the general principle that restricting ICL hurts flexibility (few-shot generalization) and improves retention.